what will bea the volume of a given mass of a gas at a pressure of 50cm of Hg if it occupies 260mL at a pressure of 98 cm of hg keeping the temperature constant

Grade:11

1 Answers

Khimraj
5 years ago
 
Use boyle's law
P1 * V1 = P2 * V2
Put the values from the queations
98* 260 = 50 * V2
V2 = 98*260/50
V2 = 509.6 ml